# More Sample-Efficient Tuning of Particle Accelerators with Bayesian Optimization and Prior Mean Models

### v2

@@ Abstract
Tuning particle accelerators is a challenging and time-consuming task, but can be automated and carried out efficiently through the use of suitable optimization algorithms. With successful applications at various facilities, Bayesian optimization using Gaussian process modeling has proven to be a particularly powerful tool to address these challenges in practice. One of its major benefits is that it allows incorporating prior information, such as knowledge about the shape of the objective function or predictions based on archived data, simulations or surrogate models, into the model. In this work, we propose the use of a neural network model as an efficient way to include prior knowledge about the objective function into the Bayesian optimization process to speed up convergence. We report results obtained in simulations and experiments using neural network priors to perform optimization of electron and heavy-ion accelerator facilities, specifically the Linac Coherent Light Source and the Argonne Tandem Linear Accelerator System. Finally, we evaluate how the accuracy of the prior mean predictions affect optimization performance.
I Introduction
@@
Particle accelerators are complex machines that have thousands of free parameters which can be tuned to increase accelerator performance and beam quality. Controlling these parameters in real time during accelerator operations through the use of advanced algorithms, so-called “online” control, enables more complex beam manipulation capabilities and reduces the need for facility operators to conduct routine tuning tasks. However, using optimization algorithms to solve facility-scale optimization problems that contain hundreds of free parameters is an open challenge due to the exponential scaling of parameter space volume with increasing numbers of parameters, often referred to as the “curse of dimensionality” [1]. Addressing facility-scale optimization problems requires control algorithms that can efficiently find solutions in extremely large parameter spaces.
Bayesian optimization (BO) [2, 3] is a model-based optimization algorithm that has shown particular promise towards solving online accelerator control problems in a variety of contexts [4]. This control algorithm builds Gaussian process models (GPs) [5] that use Bayesian statistics to predict objective function values (and corresponding uncertainties) from scratch using previous measurements of the objective function. As a result, BO algorithms can find solutions in a sample-efficient manner, which is especially important for solving problems where evaluating the objective function is expensive.
The convergence speed of BO algorithms can be improved further by leveraging its ability to easily encode prior information about the objective function into the model, instead of building the model from scratch every time the optimization is run. Adding prior information to the GP model used in BO improves the model accuracy, leading to faster convergence to optimal solutions, especially when the parameter space is larger than a few variables. In the context of accelerator physics, this type of information can be gathered from a variety of sources, including basic beam dynamics principles, previous optimization runs, historical data sets, and detailed physics simulations.
A straightforward method for incorporating prior information into the GP model is to add available measurement or simulation data directly to the model. This can also be done using multi-task GP models [6], which utilize learned correlations between prior data and measured values to improve the accuracy of model predictions. However, both of these methods suffer from the poor computational scaling of GP models (), preventing the use of large training data sets for online modeling applications.
An alternative method for incorporating physics information into the GP model is to specify a physics-informed kernel function, as demonstrated in [7, 8]. Incorporating physics information into the kernel function allowed the GP model to reduce the data needed to accurately predict cross-correlations between adjacent quadrupole focusing magnet strengths, resulting in improved convergence time when optimizing free electron laser parameters. However, this method is limited to defining correlations between function values at different points in input space, not the absolute value of the function.
In this work, we demonstrate and evaluate a method for incorporating prior physics information into GP models by combining them with fast-executing neural network (NN) surrogate models of the target systems, trained on historical data and beam dynamics simulations. We expand on previous work in [9, 10, 11] to show that combining approximate or well-calibrated NN models with GP models is an effective, computationally inexpensive way to incorporate detailed physics information into BO algorithms, significantly reducing the number of optimization steps needed to converge to a solution. We discuss suitable metrics for model accuracy and the systematic dependency of BO performance on model quality. We also developed software tools, which we describe briefly in Appendix C, to facilitate easy incorporation of prior models into the optimization package Xopt [12], making this technique more broadly accessible to the accelerator community.
We present a simulated study of how incorporating NN prior models impacts optimization of a simulated version of the Linac Coherent Light Source (LCLS) photoinjector [13]. Finally, we present experimental results from using NN prior models trained on simulation at the Linac Coherent Light Source (LCLS) and use NN models to perform transfer learning from one beam species to another at the Argonne Tandem Linear Accelerator System (ATLAS) [14].
II Bayesian Optimization
Bayesian optimization is a sample-efficient, global optimization routine for black-box functions that are typically expensive to evaluate. While this section provides a brief introduction to the basic concepts, we refer to [4] for a more detailed description of BO in the context of particle accelerators.
Assuming the observable target value at input parameter is given by
where is additive noise following a Gaussian distribution, , the problem addressed by BO is to find
with and , in as few steps as possible. To achieve its high sample efficiency BO employs a statistical surrogate model of the objective function, generally represented as a GP, which informs its search in input space. The GP is typically built from scratch during optimization. Based on the belief that the objective function is drawn from a prior probability distribution, it leverages Bayes’ theorem to find the posterior distribution based on the likelihood of the data samples collected so far. The GP surrogate model thus provides not only value predictions for , but a full predictive distribution including the uncertainties associated with those predictions.
The surrogate model is then passed to an acquisition function, which uses the model to predict the value of potential future measurements towards finding a global extremum. Popular acquisition functions, such as Expected Improvement (EI) and Upper Confidence Bound (UCB) [15], strike a balance between choosing points that add more information to the global GP model (exploration) and choosing points that are near predicted extremum (exploitation). Points in parameter space that maximize the acquisition function are then observed experimentally and added to the model data set. This process repeats sequentially until a convergence criteria is met. Regardless of the choice of acquisition function, the performance of BO algorithms depends on accurate GP surrogate modeling of the objective function using as few measurements as possible, which we examine in this work.
II.1 Standard Gaussian Process Modeling
A Gaussian process [5] is a distribution of functions denoted as
with mean and variance given by
where denotes the covariance matrice between the respective sets of data and is the identity matrix.
Note that to account for a constant mean function that is not zero, i.e., , we can simply redefine the objective function as
and still use Eq. (5) to compute the posterior mean.
III Non-Constant Prior Means
Although standard BO is generally a sample-efficient algorithm in many practical applications, more challenging optimization problems like those with a high-dimensional search space can still take prohibitively many steps to reach the desired optimization target. As the dimensionality of the problem increases, so does the amount of data needed to make accurate global predictions of the objective function. However, prior information can easily be encoded into GP models. If there is sufficient prior knowledge about the objective function to build an a-priori model, this initial estimate can serve as a more informative prior mean function for the GP and thereby improve the model’s predictive accuracy and BO convergence speed.
Incorporating a non-zero prior mean function into a GP model re-incorporates an extra term ignored in Eq. 5, producing posterior mean function values
with . For test points that are far away from previous measurements in parameter space (), the posterior mean function values are equal to the prior mean values at the test points . This effect is illustrated in Fig. 1, where the mean of the posterior distribution reverts back to the prior mean as the distance between test points and training data increases. If the prior mean function accurately predicts the objective function, the GP model can make similarly accurate predictions of the objective without any data. Conversely, if portions of the prior mean incorrectly make predictions, the posterior predictions of the GP model will reflect updated values from training data. In this way, the GP can be interpreted as a model of the difference between the prior mean function and the true objective function.
The prior mean functions can be represented in a number of different ways, including analytic functions, surrogate models or even detailed physics simulations. Ideal models for prior mean function specification are fast executing and are differentiable, i.e., they efficiently compute derivatives of function values with respect to input parameters [16]. These properties are important for fast optimization of the BO acquisition function to select future input parameters.
One such surrogate model which has these properties is a neural network (NN). Neural networks can be trained using a variety of information sources including historical experimental data, simulation data, and physics principles. Additionally, NNs are inherently differentiable owing to the need for gradient calculation for effective model training.
Finally, a critical aspect of NN models is their relatively short evaluation time, which is independent of the training data set size. This enables incorporating extremely large amounts of data into the model from historical measurements or simulations, thereby circumventing the poor computational scaling of GP models. This is demonstrated in Fig. 2, which compares the computational cost of evaluating GP models of the LCLS photoinjector (see Sec. IV) on a fixed number of points with different methods of incorporating prior data into the GP model. If we were to directly incorporate large training data sets into the GP model, the evaluation time scales as . However, if we use a neural network to represent the prior information in the data set the evaluation time of the GP model is substantially smaller and independent of the number of training points. This computational scaling makes a significant practical difference when using these algorithms for online optimization, especially at user facilities that are in high-demand and need to spend as little time as possible tuning the beam.
In practice, and in the cases described here, the NN surrogate model is trained prior to performing optimization using BO. While this does require upfront effort to construct and train a NN surrogate model, the reduction in evaluation time for optimizing high-dimensional optimization problem makes up for the additional work, especially if the optimization is performed repeatedly. Furthermore, NN surrogate models of accelerators are already generated for various applications such as virtual diagnostics [17], virtual accelerator models for online operator feedback [18], and accelerator design [19]. By representing the prior mean as a neural network, we combine the data efficiency of the GP model with the better computational scaling and ability to represent large amount of previous data provided by NNs.
IV Simulation Studies
@@
In order to better understand the influence of a non-constant prior mean on BO performance, its effects are initially studied in simulations using a surrogate model for the LCLS photoinjector. The beamline is parameterized by scalar parameters that control the laser spot size on the cathode, the laser pulse length, the bunch charge, the phases and amplitudes of two RF accelerating cavities, and the focusing strengths of a number of solenoid, normal quadrupole and skew quadrupole magnets. We use a fully connected NN surrogate model trained on data from the beam dynamics simulation code IMPACT-T [20] to predict five scalar quantities of the beam distribution (beam sizes and transverse emittances) at an optical transition radiation diagnostic screen downstream of the photoinjector (OTR2). This surrogate model is incorporated into the GP model built by Xopt [12] as a prior mean function using the LUME-Model [18] framework to perform calibration and scaling to experimental units. A detailed description of the surrogate model can be found in Appendix B, practical details of incorporating the prior mean model into the GP are discussed in Appendix C.
@@
We run BO to minimize the transverse beam size of a round beam with respect to machine parameters that control the magnet and RF settings (with the laser spot size, laser pulse length, and bunch charge parameters held fixed) using the objective function
@@
where is the Pearson correlation coefficient for a collection of samples, which is typically a more accurate metric in assessing model quality and predicting BO performance. Furthermore, as a slightly more intuitive measure of the absolute error than MSE, we augment the correlation coefficient with the mean absolute error (MAE) to describe the overall model quality. The respective values for the trained neural networks are calculated individually for each run and are summarized in Table 1.
| Epochs | Correlation | MAE () |
@@
V LCLS Demonstration
The LCLS surrogate model trained on data generated by the LCLS IMPACT-T model was then used as an NN prior mean to perform online optimization of the LCLS photoinjector beamline. We repeatedly performed BO with the EI acquisition function starting with a fixed set of three randomly chosen initial points, with and without the NN prior mean function, to minimize the objective function in Eq. 9. During this, the laser spot size, pulse length, RF cavity parameters and bunch charge were fixed, resulting in a 9-dimensional optimization problem. Results from these optimization runs can be seen in Fig. 5. We observe that using the NN prior model consistently outperforms the constant prior in the first optimization steps. It takes the constant prior five steps longer on average than the NN prior to reach the same optimum. However, after approximately ten steps, the constant prior model surpasses the optimum achieved by the NN prior model, likely due to its low correlation with the measured objective (, ). The NN prior mean model provides an advantage in speed during the early stages of coarse optimization, and it is then surpassed in later refinement of the optimum by standard BO. Based on the simulation results in Sec. IV, it is expected that NN models that are better calibrated to experimental data will increase the performance of BO in this context. In the meantime, it is possible to improve the optimization performance at later steps by transitioning from the NN model to a constant prior mean during the optimization, as discussed in Appendix A.
VI ATLAS Demonstration
The main objective of beam optimizations at ATLAS is to maximize beam transmission to the target while preserving overall beam quality. In this application, the focus is to maximize beam transmission through the ATLAS material irradiation station (AMIS) line by adjusting the current settings of a magnetic quadrupole triplet and doublet, totaling five parameters.
We began with establishing the baseline performance of standard BO with a constant prior mean, initiating the optimization with only one data point, resulting in an initial transmission close to , and employed the Upper Confidence Bound (UCB) acquisition function with . This algorithm was executed on the machine using an 16O beam times, each time starting from the same conditions. Refer to Fig. 6, where the blue dashed line represents the median maximum transmission at each step of the constant prior BO optimization, with the surrounding area depicting the confidence level.
We then trained a neural network model on a dataset comprising experimental data points obtained from a prior experiment involving a 14N heavy-ion beam to predict transmission as a function of the quadrupole configurations, details of which are discussed in Appendix B. Doing this, we demonstrate and take advantage of effective transfer learning, where historical data gathered during ATLAS operations using one species, 14N, is used to inform optimization of accelerator operations for a different but similar heavy-ion species, 16O.
The trained NN model was then incorporated into the BO algorithm as the prior mean and used for optimization with the 16O beam, employing the same initial conditions as the constant prior BO optimization mentioned earlier. This process was repeated times as well. The results, depicted in orange in Fig. 6, clearly illustrate the superior performance of BO with a neural network prior mean. This approach outperforms traditional BO methods with no prior knowledge, underscoring not only the advantage of incorporating a prior mean model in Gaussian processes but also the potential of transfer learning techniques in optimizing particle accelerators by using models trained under different beams and conditions.
However, as emphasized earlier, the accuracy of the surrogate model holds a pivotal role in determining the effectiveness of the optimization process. This critical factor was empirically validated at ATLAS through an extension of the same BO with GP framework previously mentioned, by employing surrogate models with different accuracies. These surrogate models were trained using the same dataset but with varying duration of training, yielding a spectrum of accuracies measured by different correlations between their predictions and the actual transmission in the machine. The respective values for the trained neural networks are listed in Table 2. Figure 7 illustrates how the performance of BO augmented by an NN prior mean exhibits marked improvement with the increasing correlation between the model’s predictions and the machine’s behavior.
| Epochs | Correlation | MAE () |
VII Conclusion
In this work we have demonstrated how incorporating a fast executing NN surrogate model of the target system can dramatically improve the performance of BO algorithms. In cases where the prior model is more accurate and shows stronger correlation with ground truth values, the algorithm significantly outperforms standard BO with a constant prior mean function. These NN based prior mean models can be used to efficiently incorporate information from beam dynamics simulations, historical data, and/or similar accelerator operating configurations to accelerate online tuning of accelerators, all while quickly adapting to discrepancies between surrogate model predictions and observations on the real machine. However, care must be taken if the prior mean model is poorly correlated or anti-correlated with the true objective function, which can reduce BO performance below the performance with a constant prior.
As was observed in the ATLAS demonstration, the use of an expressive prior mean model can also facilitate transfer learning across different beam setups. This could aid rapid switching between beam setups at accelerators that serve a variety of experiments with different requirements (e.g. different species of beams as in the ATLAS case, or different electron beam characteristics requested by photon science users at the LCLS).
Future work in this area should focus on studying metrics for training prior models that yield good performance in the context of BO, including disentangling the effects of the two metrics identified in this work to describe model quality, i.e., MAE and correlation. Additionally, as discussed in Appendix A, it is possible to recover optimization performance on-the-fly, i.e., during optimization, when the prior mean model is a poor predictor of the objective function; future work should more thoroughly explore different heuristics for weighting the use of a NN prior mean function and trainable constant prior functions to improve performance. Furthermore, differentiable beam dynamics models, such as those implemented in Bmad-X [22] or Cheetah [23, 24] should also start to be incorporated into GP models as physics-informed priors. Finally, non-constant prior mean functions can be incorporated into GP models with constraining functions for use in constrained optimization (see [4] for details) to reduce the number of constraint violations during optimization.
Acknowledgements.
Author contributions: Conceptualization, T.B., J.L.M., B.M., D.R., R.R., and A.L.E.; Data curation, T.B., J.L.M., C.X., and K.R.L.B.; Formal analysis, T.B. and J.L.M.; Funding acquisition, D.R., A.L.E., and B.M.; Investigation, T.B. and J.L.M.; Methodology, R.R., T.B., J.L.M., and A.L.E.; Software, R.R., C.X., K.R.L.B., T.B., and J.L.M.; Experiment, T.B., J.L.M., R.R., and A.L.E.; Supervision, D.R., A.L.E., R.R., and B.M.; Validation, T.B. and J.L.M.; Visualization, T.B. and J.L.M.; Writing—original draft, T.B., J.L.M., R.R., B.M., and A.L.E.; Writing—review and editing, T.B., J.L.M., R.R., D.R., B.M., and A.L.E. All authors have read and agreed to the published version of the manuscript. This work was funded by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. It was also supported by the U.S. Department of Energy, Office of Nuclear Physics under Contract No. DE-AC02-06CH11357. The presented research used the ATLAS facility, which is a DOE Office of Science User Facility.Appendix A Low-Quality Prior Mean Models
@@ Acknowledgements.
Appendix B Neural Network Models

### v3

# Leveraging Prior Mean Models for Faster Bayesian Optimization of Particle Accelerators

@@ Abstract
Tuning particle accelerators is a challenging and time-consuming task that can be automated and carried out efficiently using suitable optimization algorithms, such as model-based Bayesian optimization techniques. One of the major advantages of Bayesian algorithms is the ability to incorporate prior information about beam physics and historical behavior into the model used to make control decisions. In this work, we examine incorporating prior accelerator physics information into Bayesian optimization algorithms by utilizing fast executing, neural network models trained on simulated or historical datasets as prior mean functions in Gaussian process models. We show that in ideal cases, this technique substantially increases convergence speed to optimal solutions in high-dimensional tuning parameter spaces. Additionally, we demonstrate that even in non-ideal cases, where prior models of beam dynamics do not exactly match experimental conditions, the use of this technique can still enhance convergence speed. Finally, we demonstrate how these methods can be used to improve optimization in practical applications, such as transferring information gained from beam dynamics simulations to online control of the LCLS injector, and transferring knowledge gained from experimental measurements across different operating modes, such as accelerating different ion species at the ATLAS heavy ion accelerator.
Introduction
@@
Particle accelerators are complex machines that have thousands of free parameters which can be tuned to increase accelerator performance and beam quality. Controlling these parameters in real time during accelerator operations through the use of advanced algorithms, so-called online control, enables more complex beam manipulation capabilities and can help reduce the time spent on conducting routine tuning tasks. However, using optimization algorithms to solve facility-scale optimization problems that contain hundreds of free parameters is an open challenge due to the exponential scaling of parameter space volume with increasing numbers of parameters, often referred to as the “curse of dimensionality” [1]. When developing effective algorithms for accelerator control, it is important to incorporate prior information about beam physics into the algorithm to efficiently find solutions in extremely large parameter spaces. Additionally, many accelerator facilities frequently switch beam conditions for different users. This is the case, for example, at both the Linac Coherent Light Source (LCLS) and the Argonne Tandem Linear Accelerator System (ATLAS), a heavy ion accelerator. While these configurations can be treated as separate optimization problems, the time needed to switch between conditions could be reduced by leveraging information from related machine configurations.
Bayesian optimization (BO) [2] is one such algorithm type that has been used in accelerators and can incorporate prior physics information into the decision-making process. The BO algorithm builds a Gaussian process model (GP) [3] that uses Bayesian statistics to predict objective function values (and corresponding uncertainties) from measurements of the objective function. This model is then used to inform control decisions in search of a globally optimal parameter set. Previous work [4, 5, 6] has demonstrated that including prior information from physics into BO algorithms increases their ability to efficiently converge to optimal solutions in high-dimensional parameter spaces. The use of Bayesian statistics (which inherently takes prior information into account) to build the GP model provides a natural framework for incorporating physics information into the modeling of accelerator-related optimization objectives. For example, GP models are built using a prior mean function, which specifies the prediction of the GP model in the absence of data. These prior mean functions are usually specified to be uninformative, as shown in Fig. 1(a). However, if prior information about the objective function is known, either from beam dynamics laws, simulations or prior data, an informative prior mean can be specified to incorporate this information directly into the GP model, as shown in Fig. 1(b). This can increase the accuracy of the GP model which helps guide the search for an optimum, improving the algorithm’s ability to converge to solutions quickly in high-dimensional spaces.
In this work, we demonstrate and evaluate the use of fast-executing neural network (NN) surrogate models trained on beam dynamics simulations and historical data sets, as informative, physics-based prior mean functions in GP models to improve online accelerator control for two impactful use cases: injector tuning for the LCLS and tuning for different ion species at ATLAS. In particular, we demonstrate applications of this method in the context of realistic accelerator operations, including adapting surrogate models trained on beam dynamics simulations to online control at the LCLS injector, and transfer learning between different ion species at the ATLAS heavy ion accelerator, with improvements in convergence speed. These two cases were carefully chosen based on the expected benefit of the NN prior approach. The LCLS has to frequently switch setups for different users (different beam energies, bunch lengths, etc), and the injector is one of the more challenging parts of the machine to optimize due to space charge forces and the resulting sensitivity of the system to changes in controlled and uncontrolled inputs (such as drift in the initial laser spot). The injector is also not continuously tuned, because the relevant diagnostics are destructive. This makes it appealing both for using BO and for leveraging NN priors obtained from simulations and historical data. In addition, detailed physics simulations and ML-based surrogates of the LCLS injector are readily available. Similarly, ATLAS has to switch frequently between different ion species. Complicating this, beam dynamics for heavy ion linacs are very complicated and sensitive to machine inputs and changing conditions. Leveraging information across different configurations to speed up switching between ion species would substantially aid operation. The work presented here expands on previous work [7, 8, 9] that was conducted in a more limited scope. The first shows an offline proof-of-concept demonstration of using a NN prior for emittance and matching tuning on the LCLS injector. The second is a preliminary experimental demonstration at FRIB using an NN prior from historical data for tuning. The third is an initial experimental demonstration at ATLAS. We also explore the impact that the accuracy of the prior model has on the speed of convergence, and present a method to counteract the effects of an inaccurate model. As physics models often deviate substantially from real machine behavior, handling of machine-model mismatch is a major consideration for practical use of the NN prior approach. We show that combining approximate or well-calibrated NN models with GP models is an effective, computationally inexpensive way to incorporate detailed physics information into BO algorithms, significantly reducing the number of optimization steps needed to converge to a solution, even in cases where the NN model does not perfectly predict the experimental objective. Our demonstration also shows the utility of using prior information across ion species to improve tuning speed, which is important for current and future nuclear physics machines.
Results
Here we present simulated and experimental studies, comparing the optimization performance of BO algorithms with and without using NN models as prior means in GP models.
Simulation Studies
@@ Results
In order to understand the influence of a non-constant prior mean on BO performance, its effects are initially studied in simulations using a surrogate model for the LCLS photoinjector, the layout of which is shown in Fig.2. The beamline is parameterized by 16 scalar parameters that control the laser spot size on the cathode, the laser pulse length, the bunch charge, the phases and amplitudes of two RF accelerating cavities, and the focusing strengths of a number of solenoid, normal quadrupole and skew quadrupole magnets. We use a fully connected NN surrogate model trained on data from the beam dynamics simulation code IMPACT-T [10] to predict five scalar quantities of the beam distribution (beam sizes and transverse emittances) at an optical transition radiation diagnostic screen downstream of the photoinjector (OTR2). This surrogate model is incorporated into the GP model built by Xopt [11] as a prior mean function using the LUME-Model [12] framework to perform calibration and scaling to experimental units.
@@
We run BO to minimize the transverse beam size of a round beam with respect to of the above machine parameters that control the magnet and RF settings (with the laser spot size, laser pulse length, and bunch charge parameters held fixed) using the objective function
@@
GP Prediction Scaling with NN Priors
A critical advantage of NN models as prior mean functions is their relatively short evaluation time, which is independent of the training data set size. This enables incorporating extremely large amounts of data into the model from historical measurements or simulations, thereby circumventing the poor computational scaling of GP models. This is demonstrated in Fig. 4, which compares the computational cost of evaluating GP models on a fixed number of points with different methods of incorporating prior data into the GP model. If we were to directly incorporate large training data sets into the GP model, the evaluation time scales as . However, if we use a neural network to represent the prior information in the data set the evaluation time of the GP model is substantially smaller and independent of the number of training points. This computational scaling makes a significant practical difference when using these algorithms for online optimization, especially at user facilities that are in high-demand and need to spend as little time as possible tuning the beam.
Performance Recovery for Low-Quality Prior Mean Models
@@
LCLS Experimental Demonstration
The LCLS surrogate model trained on data generated by the LCLS IMPACT-T model was used as an NN prior mean to perform online optimization of the LCLS photoinjector beamline. We repeatedly performed BO with the EI acquisition function starting with a fixed set of three randomly chosen initial points, with and without the NN prior mean function, to minimize the objective function in Eq. 1. During this, the laser spot size, pulse length, RF cavity parameters and bunch charge were fixed, resulting in a 9-dimensional optimization problem. Results from these optimization runs can be seen in Fig. 6. We observe that using the NN prior model consistently outperforms the constant prior in the first optimization steps. It takes the constant prior version five steps longer on average than the NN prior version to reach the same optimum. However, after approximately ten steps, the constant prior model surpasses the optimum achieved by the NN prior model, likely due to its low correlation with the measured objective (, ). The NN prior mean model provides an advantage in speed during the early stages of coarse optimization, and it is then surpassed in later refinement of the optimum by standard BO. Based on the simulation results discussed above, it is expected that NN models that are better calibrated to experimental data will increase the performance of BO in this context. In the meantime, it is possible to improve the optimization performance at later steps by transitioning from the NN model to a constant prior mean during the optimization, as demonstrated in the simulation section.
ATLAS Experimental Demonstration
The main objective of beam optimizations at ATLAS is to maximize beam transmission to the target while preserving overall beam quality. In this application, the focus is to maximize beam transmission through the ATLAS material irradiation station (AMIS) line, shown in Fig. 7 by adjusting the current settings of a magnetic quadrupole triplet and doublet, totaling five parameters. We began with establishing the baseline performance of standard BO with a constant prior mean, initiating the optimization with only one data point, resulting in an initial transmission close to , and employed the Upper Confidence Bound (UCB) acquisition function with . This algorithm was executed on the machine using an 16O beam times, each time starting from the same conditions. Refer to Fig. 8, where the blue dashed line represents the median maximum transmission at each step of the constant prior BO optimization, with the surrounding area depicting the confidence level. We then trained a neural network model on a dataset comprising experimental data points obtained from a prior experiment involving a 14N heavy-ion beam to predict transmission as a function of the quadrupole configurations. Doing this, we demonstrate and take advantage of effective transfer learning, where historical data gathered during ATLAS operations using one species, 14N, is used to inform optimization of accelerator operations for a different but similar heavy-ion species, 16O. The trained NN model was then incorporated into the BO algorithm as the prior mean and used for optimization with the 16O beam, employing the same initial conditions as the constant prior BO optimization mentioned earlier. This process was repeated times as well. The results, depicted in orange in Fig. 8, clearly illustrate the superior performance of BO with a neural network prior mean. This approach outperforms traditional BO methods with no prior knowledge, underscoring not only the advantage of incorporating a prior mean model in Gaussian processes but also the potential of transfer learning techniques in optimizing particle accelerators by using models trained under different beams and conditions.
However, as emphasized earlier, the accuracy of the surrogate model holds a pivotal role in determining the effectiveness of the optimization process. This critical factor was empirically validated at ATLAS through an extension of the same BO with GP framework previously mentioned, by employing surrogate models with different accuracies. These surrogate models were trained using the same dataset but with varying duration of training, yielding a spectrum of accuracies measured by different correlations between their predictions and the actual transmission in the machine. Figure 9 illustrates how the performance of BO augmented by an NN prior mean exhibits marked improvement with the increasing correlation between the model’s predictions and the machine’s behavior.
Discussion
In this work we have experimentally demonstrated how incorporating a fast executing NN surrogate model of the target system can dramatically improve the performance of BO algorithms for two important use-cases: injector optimization at LCLS and switching between ion species at the ATLAS heavy ion linac. We also show in a simulation example that in cases where the prior model is more accurate and shows stronger correlation with ground truth values, the algorithm significantly outperforms standard BO with a constant prior mean function. These NN based prior mean models can be used to efficiently incorporate information from beam dynamics simulations, historical data, and similar accelerator operating configurations to accelerate online tuning of accelerators, all while quickly adapting to discrepancies between surrogate model predictions and observations on the real machine. However, care must be taken if the prior mean model is poorly correlated or anti-correlated with the true objective function, which can reduce BO performance below the performance with a constant prior. As was observed in the ATLAS experimental demonstration, the use of an expressive prior mean model can also facilitate transfer learning across different beam setups. This can aid rapid switching between beam setups at accelerators that serve a variety of experiments with different requirements (e.g. different species of beams as in the ATLAS case, or different electron beam characteristics requested by photon science users at the LCLS).
Future work in this area will focus on studying metrics for training prior models that yield good performance in the context of BO, including disentangling the effects of the two metrics identified in this work to describe model quality, i.e., MAE and correlation. Additionally, as discussed in the simulation section, it is possible to recover optimization performance on-the-fly, i.e., during optimization, when the prior mean model is a poor predictor of the objective function; future work should more thoroughly explore different heuristics for weighting the use of a NN prior mean function and trainable constant prior functions to improve performance. Furthermore, differentiable beam dynamics models, such as those implemented in Bmad-X [14] or Cheetah [15, 16] should also start to be incorporated into GP models as physics-informed priors. Finally, non-constant prior mean functions can be incorporated into GP models with constraining functions for use in constrained optimization (see [2] for details) to reduce the number of constraint violations during optimization.
Methods
We begin with a discussion of GP modeling and the method by which we incorporate prior mean functions into the modeling process. We then discuss how we characterized prior model quality, construction, and software implementation details regarding how NN models were integrated into GP models using the software package Xopt [11].
Gaussian Process Modeling and Bayesian Optimization
A Gaussian process [3] (GP) is a distribution of functions denoted as
with mean and variance given by
where denotes the covariance matrix between the respective sets of data and is the identity matrix. Note that to account for a constant mean function that is not zero, i.e., , we can simply redefine the objective function as
and still use Eq. (6) to compute the posterior mean.
The GP model is then passed to an acquisition function, which uses the model to predict the value of potential future measurements towards finding a global extremum. Popular acquisition functions, such as Expected Improvement (EI) and Upper Confidence Bound (UCB) [17], strike a balance between choosing points that add more information to the global GP model (exploration) and choosing points that are near predicted extremum (exploitation). Points in parameter space that maximize the acquisition function are then observed experimentally and added to the model data set. This process repeats sequentially until a convergence criteria is met. Regardless of the choice of acquisition function, the performance of BO algorithms depends on accurate GP surrogate modeling of the objective function using as few measurements as possible, which we examine in this work.
Non-Constant Prior Means
If there is sufficient prior knowledge about the objective function to build an a-priori model, this initial estimate can serve as a more informative prior mean function for the GP and thereby improve the model’s predictive accuracy and BO convergence speed. Incorporating a non-zero prior mean function into a GP model re-incorporates an extra term ignored in Eq. 6, producing posterior mean function values
with . For test points that are far away from previous measurements in parameter space (), the posterior mean function values are equal to the prior mean values at the test points .
This effect is illustrated in Fig. 10, where the mean of the posterior distribution reverts back to the prior mean as the distance between test points and training data increases. If the prior mean function accurately predicts the objective function, the GP model can make similarly accurate predictions of the objective without any data. Conversely, if portions of the prior mean incorrectly make predictions, the posterior predictions of the GP model will reflect updated values from training data. In this way, the GP can be interpreted as a model of the difference between the prior mean function and the true objective function.
Metrics for Characterizing Model Accuracy
where is the Pearson correlation coefficient for a collection of samples, which is typically a more accurate metric in assessing model quality and predicting BO performance. Furthermore, as a slightly more intuitive measure of the absolute error than MSE, we augment the correlation coefficient with the mean absolute error (MAE) to describe the overall model quality.
Neural Network Models